# Generalized Landau Yang Theorem

### v1

@@ Abstract
Landau Yang theorem is well known for the past several decades. It prohibits decay of a massive spin 1 particle to two photons. This emerges simply from the representation theory of the Poincare group and Bose Statistics. It does not require any action or Lagrangian. We generalize this theorem to theories with supersymmetry (SUSY) which disallows even decay to two photinos (Majorana fermions). We will prove that if the photon has a mass, howsoever small, this theorem can be evaded. We also show that the supersymmetric selection rule above can also be evaded through the Stueckelberg mass term. Further interesting implications are also pointed out.
Generalized Landau Yang Theorem
T. R. Govindarajan111trg@imsc.res.in, govindarajan.thupil@krea.edu.in
The Institute of Mathematical Sciences, Chennai 600113, Tamil Nadu, India
Krea University, Sricity, 517646, Andhra Pradesh, India
Rakesh Tibrewala222rtibs@lnmiit.ac.in
The LNMIIT, Jaipur 302031, Rajasthan, India
October 07, 2024
1 Introduction:
Landau[1] and Yang [2] independently proved a theorem which disallows a massive spin 1 particle from decaying to two massless spin-1 bosons without further internal symmetries. The elegance of the theorem stems from the fact that it does not depend on any interactions and is very generic. This is easy to understand from the representation theory of the Poincare group and Bose-Einstein statistics. For example, the $\rho$ meson which has spin 1 does not decay to two photons whereas the pion (which has spin 0) does. It prohibits the eletroweak $Z$ particle to decay to two photons (the bound on the branching ratio $\Gamma_{Z\rightarrow\gamma\gamma}/\Gamma_{\mathrm{Tot}}<5.2\times 10^{-5}$ [3]). Recently this theorem has been applied to understand that the Higgs Boson cannot be a spin 1 particle due to its decay to two photons [4, 5].
@@ 1.1 Simple proof
Massive spin 1 particle can be taken to be at rest with the four momentum $P^{\mu}~{}=~{}(m,0,0,0)$. Its spin can be given as $\vec{S}$. The massless photons will have momenta $\vec{k}$ and $-\vec{k}$ in order to conserve the momentum. The polarization of these two photons will be in the plane perpendicular to $\vec{k}$. We can call them as $\vec{s}_{1},\vec{s}_{2}$. They will obey
@@
${\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})~{}=~{}\sum_{1}^{3}a_{i}(k^{2})~{}A_{i}$ (2)
Here $a_{i}(k^{2})$ are arbitrary invariant functions. Now the Bose symmetry of photons implies that the amplitude should be the same if we transform $\vec{k}\rightarrow-\vec{k}$ and interchange $\vec{s}_{1}~{}\leftrightarrow~{}\vec{s}_{2}$. That is we should have
@@ 1.2 Representation theory of Poincare symmetry
We recollect [6] that massless irreducible representations of the Poincare group are given by the helicity. On the other hand, massive representations are specified by the mass parameter as well as the spin value coming from the Pauli-Lubanski vector. We also need to represent the massless photons covariantly using a four vector $A_{\mu}$ to discuss further interactions through gauge theory. That brings in the helicity ±$\pm~{}1$. These follow from the induced representations of the little group of four momentum vector $p^{\mu}~{}=~{}(\omega,0,0,\omega)$ of a massless particle which is the Euclidean group in two dimensions. The little group for a massive particle on the other hand is $SO(3)$ which comes from the fact that the massive particle can be brought to rest in a frame and its four momentum in that frame is $(m,0,0,0)$. In addition it is well known that the helcity ±$\pm~{}1$ representations from $A_{\mu}$ are reducible but not decomposable[7]. We need to look at the content of the product of two such helicity representations through Clebsch Gordon coefficients. Also Bose symmetry should be imposed. It is clear that the decay of a massive spin 1 particle to two photons to be possible demands that the symmetric product of two massless representations of the Poincare group contains a massive spin one representation. It was shown by Balachandran and Jo[8] that it does not contain such a representation. The direct product of two massless representations with four momenta $(p,0,0,\pm p)$ and helicities $\lambda_{1},\lambda_{2}$ can be given by:
@@
where $d\xi(R)$ is the Haar measure on $SU(2)$ and $D_{\mu~{}\lambda_{1}-\lambda_{2}}^{j}$ are the spin-j representation matrices. For the case of two identical particles we have
$|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}=~{}\frac{1}{2}\left(|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\pm~{}(-1)^{j+\lambda_{1}+\lambda_{2}}~{}|\lambda_{2}~{}\lambda_{1}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\right)$ (6)
For photons obeying Bose statistics we need the symmetric product. For the symmetric product with two identical particles of helicities $\lambda_{i}=\pm 1,i=1,2$ we have for $\lambda_{1}=\pm\lambda_{2}$:
@@
With $j=1$ we find the right hand side vanishes. This implies two photon states cannot have $j=1$ component thereby forbidding, for example, $Z\rightarrow 2\gamma$ transition and many others. On the other hand if $j=0$ as we expect for the Higgs, the decay to two photons is not prohibited by the theorem and provides a strong evidence for the scalar nature of Higgs [4, 5].
@@ 2 Generalization to SUSY
Super Poincare algebra also has two Casimir operators. The first is the usual $P_{\mu}~{}P^{\mu}$ corresponding to the mass of the particles. But the spin from the Pauli Lubanski vector does not provide the other invariant. There is a generalization of this, known as the superspin operator which commutes with all the generators of the super Poincare algebra [9]. The mass $m$ and the superspin operator $Y$ provide the massive irreducible representations. When Y = 0 we get a supersymmetric multiplet consisting of a scalar and a spin $1/2$ particle of the same mass.
@@
Hence combined with supersymmetry we will also have in addition zero amplitude for Z decaying to two photinos. Here the Fermi symmetry of photinos along with SUSY forbids such a process. Interestingly such a decay is prohibited by simply Poincare group representation theory itself – a massive spin 1 particle cannot decay to two (identical) massless Majorana particles, Fig. 1(a). This is seen in Balachandran and Jo paper through the decomposition of the antisymmetric product of two spin $1/2$ Majorana representations. To see this consider Eq. (7) with $j=1,\lambda_{i}=1/2$ corresponding to decay to two photinos. It is easy to see the right hand side of Eq.7 is identically zero.
@@
We can extend this even further without much ado to the zero amplitude for a zino going to a photon and a photino, Fig. 1(b). This is inspite of the fact that all three particles are distinct. Only representation theory of the super Poincare algebra along with symmetric/antisymmetric product of irreducible reprsentations provides such a powerful result.
3 Evading the theorem:
In the literature several options have been discussed for evading the theorem through relaxing the conditions for the theorem to hold. For example, one can consider a background of homogeneous electric or magnetic fields [10] there by relaxing the requirement of Lorentz invariance/Poincare symmetry. The external field gives another vector in addition to the polarization vectors and which therefore allows nonzero amplitude for the process. Such background magnetic fields are present in astrophysical contexts. If the massless particles are gluons then there are additional internal color degree of freedom which gives scope for decay of $Z\rightarrow g~{}g$ in QCD[11] (but the caveat is that gluons are always confined).
Much more recently, using noncommutative geometry which is considered as violating Lorentz symmetry, the possibility of $Z\rightarrow\gamma+\gamma$ has been considered. It is well known that massless nature of photon leads to infrared divergences in QED. This was studied extensively in the sixties and was related to the difficulties associated with gauge invariant description of electron which is always surrounded by soft photons (photons of infinitesimal frequencies). This led to a new possibility to evade the Landau-Yang theorem arising from the fact that due to the asymptotic properties of QED the contribution of soft photons can be included in a way which provides Lorentz noninvariant sectors [12].
@@ 4 Massive photons and Landau-Yang theorem
Unlike the above cases we consider an alternative which does not require breaking of Lorentz invariance. We will consider the photons to be massive (note that experimentally there is only an upper bound on the mass of the photon, see subsection 4.1 below). It is obvious that if the photons have a small mass the question of the representation theory of the Poincare group takes a new angle. It becomes a question of whether in the tensor product of two massive spin 1 photons obeying Bose symmetry there is a massive spin 1 representation (Z). The answer is obviously yes. This possibility arises because massive photons have three degrees of freedom including the longitudinal photons and Eq.1 is no longer valid. It gives a longitudinal component to the polarization $\vec{s}_{i}$ in the direction of $\vec{k}$ so that $k\cdot\vec{s}_{i}\neq 0$. This allows for a nonzero amplitude for the decay of a massive spin 1 particle like $Z$ to two massive $\gamma$ with a subtantial contribution from the longitudinal component for one of the photons.
@@ 4.1 Inonu-Wigner contraction
We mentioned that if the photon has a mass, even if it is small, the Landau-Yang theorem can be evaded. For this one needs to understand the nature of the limit $m_{\gamma}~{}\rightarrow 0$. The question of mass of the photon was raised by Schrodinger himself in 1955 [13] and answered by him that QED results will not be affected by a small photon mass as long as its interaction with matter is only through a conserved current, that is, if $L_{int}~{}=~{}j_{\mu}~{}A^{\mu}$ where $\partial_{\mu}~{}j^{\mu}~{}=~{}0$. Schrodinger himself worked out the upper bound of the mass of the photon using geomagnetic data. The current limit for the mass of the photon as given by the Particle data book [14] is $m_{\gamma}~{}<~{}10^{-18}eV$. The question of mass going to zero limit can be posed as group theoretic question of contraction of $SO(3)~{}\rightarrow E(2)$. The contraction of the massive representation of the Poincare group to the massless representation is discussed in Kim and Wigner [15] and here we briefly summarize the parts relevant for us. The generators of $SO(3)$ can be written as $L_{\pm},L_{3}$ with the Lie-algebra
$\left[L_{3},L_{\pm}\right]~{}=~{}\pm i~{}L_{\pm},\qquad[L_{\pm},L_{\mp}]~{}=~{}2L_{3}$ (8)
@@
$P_{x,y}~{}=~{}\pm S^{-1}L_{x,y}~{}S~{}\qquad S~{}=~{}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&R\end{pmatrix}$ (10)
@@
This was later escalated to the Poincare group by Kim and Wigner[15]. Both $SO(3)$ and $E(2)$ are subgroups of the Lorentz group and we should arrive at this contraction as part of $SL(2,C)$. This is explained better by looking at the rotation and boost generators $J_{i},~{}K_{i}$ and considering
@@
$S(R)~{}=~{}e^{-ilog(R)K_{3}}~{}=~{}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&R&0\\ 0&0&0&\frac{1}{R}\end{pmatrix}$ (12)
where $R~{}=~{}\sqrt{(1+\beta)\over(1-\beta)}$ and $\beta$ is the velocity. The generators $N_{i}$ are proportinal to $N_{i}~{}\propto~{}\frac{1}{R}\epsilon_{ij}~{}S~{}J_{j}~{}S^{-1}$
@@ 5 The process: Stueckelberg Weinberg-Salam model
It is well known that simply adding a mass term to the gauge boson Lagrangian violates the gauge invariance of the theory. To circumvent this we will consider the Stueckelberg extension of the abelian gauge boson in the Weinberg-Salam model [16, 17]. As an explicit example we consider the $Z\rightarrow 2\gamma$ process. $Z$ boson being chargeless is not coupled directly to the photon. We have to look for loop diagrams. Z interacts with the electron and the electron is coupled to the photon. The one loop diagram Fig. 2 is an example of two photon emission from $Z$ (there will be another diagram with $(p,\nu)\leftrightarrow(k,\lambda)$ which we do not show explicitly). In principle we will have contributions from the muon ($\mu$) and tau ($\tau$) loops as well as quark loops. Their contributions are expected to be small.
@@
$A_{4}~{}=~{}\Bigl{(}\vec{s}_{1}\times\vec{s}_{2}\Bigr{)}\cdot\Bigl{(}\vec{S}\times\vec{k}\Bigr{)}$ (13)
@@
$\displaystyle T$$\approx$ $\displaystyle(-1)\int\frac{d^{4}\ell}{(2\pi)^{4}}Tr\left[\gamma^{\mu}\frac{(1-\gamma^{5})}{2}\frac{i}{C_{1}}\gamma^{\lambda}\frac{i}{C_{2}}\gamma^{\nu}\frac{i}{C_{3}}\right]$ (14)
$\displaystyle+(p,\nu)\leftrightarrow(k,\lambda)$
@@
This has to be contracted with the massive gauge boson polarization vectors. As has been pointed out in [17], the Stueckelberg extension of the standard model is anomaly free and therefore we can simply estimate the amplitude for the process without worrying about the details of evaluating the loop integral. Including the coupling constants and the masses of the particles involved, we can estimate $|T|^{2}$ (upto numerical factors) to go as
@@
In the above expression $v^{Z}_{e}$ and $v^{A}_{e}$ are the effective couplings of the electron with the $Z$ boson and the photon (in the notation of [17]), $M_{Z}$ is the mass of the $Z$ boson, $m_{e}$ is the mass of the electron and $m_{\gamma}$ is the mass of the photon  [^3] Here we have projected the $Z-e-e$ vertex to the $V-A$ part since we are only interested in the order of magnitude estimate for the process. However, we note that in the Stueckelberg extension of the standard model there is a slight difference in the electromagnetic coupling of the photon to the left-handed and the right-handed electrons.. We see that in the limit $m_{\gamma}\rightarrow 0$ the $Z\rightarrow\gamma\gamma$ does not occur as required by the Landau-Yang theorem.

### v2

# Chennai 600113, Tamil Nadu, India

@@ Abstract
Landau Yang theorem is well known for the past several decades. It prohibits the decay of a massive spin 1 particle to two photons. This emerges simply from the representation theory of the Poincare group and Bose Statistics. It does not require any action or Lagrangian. We generalize this theorem to theories with supersymmetry (SUSY) which disallows even decay to two photinos (Majorana fermions) as well as the decay of a zino to a photon and a photino. We will prove that if the photon has a mass, howsoever small, this theorem can be evaded. We also show that the supersymmetric selection rule above can also be evaded through the Stueckelberg mass term. Further interesting implications are also pointed out.
1 Introduction
Landau landau and Yang yang independently proved a theorem which disallows a massive spin 1 particle from decaying to two massless spin-1 bosons without further internal symmetries. The elegance of the theorem stems from the fact that it does not depend on any interactions and is very generic. This is easy to understand from the representation theory of the Poincare group and Bose-Einstein statistics. For example, the $\rho$ meson which has spin 1 does not decay to two photons whereas the pion (which has spin 0) does. It prohibits the eletroweak $Z$ particle to decay to two photons (the bound on the branching ratio $\Gamma_{Z\rightarrow\gamma\gamma}/\Gamma_{\mathrm{Tot}}<5.2\times 10^{-5}$ PDG-z2g ). Recently this theorem has been applied to understand that the Higgs Boson cannot be a spin 1 particle due to its decay to two photons Atlas ; CMS .
@@ 1.1 Simple proof
Massive spin 1 particle can be taken to be at rest with the four momentum $P^{\mu}~{}=~{}(m,0,0,0)$. Its spin can be given as $\vec{S}$. The massless photons will have momenta $\vec{k}$ and $-\vec{k}$ in order to conserve the momentum. The polarization of these two photons, denoted $\vec{s}_{1},\vec{s}_{2}$, will be in the plane perpendicular to $\vec{k}$ so that
@@
${\cal{A}}(\vec{k},\vec{s}_{1},\vec{s}_{2},\vec{S})~{}=~{}\sum_{1}^{3}A_{i}$ (2)
Now the Bose symmetry of photons implies that the amplitude should be the same if we transform $\vec{k}\rightarrow-\vec{k}$ and interchange $\vec{s}_{1}~{}\leftrightarrow~{}\vec{s}_{2}$. That is we should have
@@ 1.2 Representation theory of Poincare symmetry
We recollect wigner that massless irreducible representations of the Poincare group are given by the helicity. On the other hand, massive representations are specified by the mass parameter as well as the spin value coming from the Pauli-Lubanski vector. We also need to represent the massless photons covariantly using a four vector $A_{\mu}$ to discuss further interactions through gauge theory. That brings in the helicity ±$\pm~{}1$. These follow from the induced representations of the little group of four momentum vector $p^{\mu}~{}=~{}(\omega,0,0,\omega)$ of a massless particle which is the Euclidean group in two dimensions. The little group for a massive particle on the other hand is $SO(3)$ which comes from the fact that the massive particle can be brought to rest in a frame and its four momentum in that frame is $(m,0,0,0)$. In addition it is well known that the helcity ±$\pm~{}1$ representations from $A_{\mu}$ are reducible but not decomposable mathews . We need to look at the content of the product of two such helicity representations through Clebsch Gordon coefficients. Also Bose symmetry should be imposed. It is clear that for a massive spin 1 particle to decay to two photons demands that the symmetric product of two massless representations of the Poincare group contains a massive spin one representation. It was shown by Balachandran and Jo bal that it does not contain such a representation. The direct product of two massless representations with four momenta $(p,0,0,\pm p)$ and helicities $\lambda_{1},\lambda_{2}$ can be given by:
@@
where $d\xi(R)$ is the Haar measure on $SU(2)$ and $D_{\mu~{}\lambda_{1}-\lambda_{2}}^{j}$ are the spin-j representation matrices. For the case of two identical particles we have the (anti)symmetric combinations
$|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}_{s,a}~{}=~{}\frac{1}{2}\left(|\lambda_{1}~{}\lambda_{2}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\pm~{}(-1)^{j+\lambda_{1}+\lambda_{2}}~{}|\lambda_{2}~{}\lambda_{1}~{}\hat{p}~{}j~{}\mu{\rangle}~{}\right)$ (6)
where the subscripts $s,a$ correspond to the upper and lower signs on the rhs respectively. For photons obeying Bose statistics we need the symmetric product. For the symmetric product with two identical particles of helicities $\lambda_{i}=\pm 1,i=1,2$ we have for $\lambda_{1}=\pm\lambda_{2}$:
@@
With $j=1$ we find the right hand side vanishes. This implies two photon states cannot have $j=1$ component thereby forbidding, for example, $Z\rightarrow 2\gamma$ transition and many others (in choi the helicity formalism is used to generalize the theorem for the decay to two identical massless particles of any spin). On the other hand if $j=0$ as we expect for the Higgs, the decay to two photons is not prohibited by the theorem and provides a strong evidence for the scalar nature of Higgs Atlas ; CMS .
@@ 2 Generalization to SUSY
Super Poincare algebra also has two Casimir operators. The first is the usual $P_{\mu}~{}P^{\mu}$ corresponding to the mass of the particles. But the spin from the Pauli Lubanski vector does not provide the other invariant. There is a generalization of this, known as the superspin operator which commutes with all the generators of the super Poincare algebra buchbinder . The mass $m$ and the superspin operator $Y$ provide the massive irreducible representations. When Y = 0 we get a supersymmetric multiplet consisting of a scalar and a spin $1/2$ particle of the same mass.
@@
Hence combined with supersymmetry we will also have in addition zero amplitude for Z decaying to two photinos. Here the Fermi symmetry of photinos along with SUSY forbids such a process. Interestingly such a decay is prohibited by simply Poincare group representation theory itself without requiring SUSY – a massive spin 1 particle cannot decay to two (identical) massless Majorana particles, figure 1(a). This was shown in Bal-Book and is seen in Balachandran and Jo paper through the decomposition of the antisymmetric product of two spin $1/2$ Majorana representations. To see this consider eq. (7) with $j=1,\lambda_{i}=1/2$ corresponding to decay to two photinos. It is easy to see that the right hand side of eq.7 is identically zero.
@@
A more interesting conclusion and which explicitly requires supersymmetry is the zero amplitude for a zino going to a photon and a photino, figure 1(b). This is inspite of the fact that all three particles are distinct. Only representation theory of the super Poincare algebra along with symmetric/antisymmetric product of irreducible reprsentations provides such a powerful result.
3 Evading the theorem
In the literature several options have been discussed for evading the theorem through relaxing the conditions for the theorem to hold. For example, one can consider a background of homogeneous electric or magnetic fields rindani there by relaxing the requirement of Lorentz invariance/Poincare symmetry. The external field gives another vector in addition to the polarization vectors and which therefore allows nonzero amplitude for the process. Such background magnetic fields are present in astrophysical contexts. If the massless particles are gluons then there are additional internal color degree of freedom which gives scope for decay of $Z\rightarrow g~{}g$ in QCD qcd ; beenakker (but the caveat is that gluons are always confined).
Much more recently, using noncommutative geometry which is considered as violating Lorentz symmetry, the possibility of $Z\rightarrow\gamma+\gamma$ has been considered behr . It is well known that massless nature of photon leads to infrared divergences in QED. This was studied extensively in the sixties and was related to the difficulties associated with gauge invariant description of electron which is always surrounded by soft photons (photons of infinitesimal frequencies). This led to a new possibility to evade the Landau-Yang theorem arising from the fact that due to the asymptotic properties of QED the contribution of soft photons can be included in a way which provides Lorentz noninvariant sectors balnew .
@@ 4 Massive photons and Landau-Yang theorem
Unlike some of the above cases, we consider an alternative which does not require breaking of Lorentz invariance. We will consider the photons to be massive (note that experimentally there is only an upper bound on the mass of the photon, see subsection 4.1 below). It is obvious that if the photons have a small mass the question of the representation theory of the Poincare group takes a new angle. It becomes a question of whether in the tensor product of two massive spin 1 photons obeying Bose symmetry there is a massive spin 1 representation (Z). The answer is obviously yes. This possibility arises because massive photons have three degrees of freedom including the longitudinal photons and eq.1 is no longer valid. It gives a longitudinal component to the polarization $\vec{s}_{i}$ in the direction of $\vec{k}$ so that $k\cdot\vec{s}_{i}\neq 0$. This allows for a nonzero amplitude for the decay of a massive spin 1 particle like $Z$ to two massive $\gamma$ with a subtantial contribution from the longitudinal component for one of the photons.
@@ 4.1 Inonu-Wigner contraction
We mentioned that if the photon has a mass, even if it is small, the Landau-Yang theorem can be evaded. For this one needs to understand the nature of the limit $m_{\gamma}~{}\rightarrow 0$. The question of mass of the photon was raised by Schrodinger himself in 1955 schrodinger and answered by him that QED results will not be affected by a small photon mass as long as its interaction with matter is only through a conserved current, that is, if $L_{int}~{}=~{}j_{\mu}~{}A^{\mu}$ where $\partial_{\mu}~{}j^{\mu}~{}=~{}0$. Schrodinger himself worked out the upper bound of the mass of the photon using geomagnetic data. The current limit for the mass of the photon as given by the Particle data book PDG is $m_{\gamma}~{}<~{}10^{-18}eV$. The question of mass going to zero limit can be posed as group theoretic question of contraction of $SO(3)~{}\rightarrow E(2)$. The contraction of the massive representation of the Poincare group to the massless representation is discussed in Kim and Wigner kim and here we briefly summarize the parts relevant for us. The generators of $SO(3)$ can be written as $L_{\pm},L_{3}$ with the Lie-algebra
$\left[L_{3},L_{\pm}\right]~{}=~{}\pm L_{\pm},\qquad[L_{+},L_{-}]~{}=~{}2L_{3}$ (8)
@@
$P_{x,y}~{}=~{}\pm\frac{1}{R}S^{-1}L_{x,y}~{}S~{}\qquad S~{}=~{}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&R\end{pmatrix}$ (10)
@@
Both $SO(3)$ and $E(2)$ are subgroups of the Lorentz group and we should arrive at this contraction as part of $SL(2,C)$ kim . This is explained better by looking at the rotation and boost generators $J_{i},~{}K_{i}$ and considering
@@
$S(R)~{}=~{}\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&R&0\\ 0&0&0&\frac{1}{R}\end{pmatrix}$ (12)
where $R~{}=~{}\sqrt{(1+\beta)\over(1-\beta)}$ and $\beta$ is the velocity and the generators $N_{i}~{}=~{}-\frac{1}{R}\epsilon_{ij}~{}S~{}J_{j}~{}S^{-1}$ where the $J_{i}$ are the angular momentum operators in the light cone coordinates.
@@ 5 The process: Stueckelberg Weinberg-Salam model
It is well known that simply adding a mass term to the gauge boson Lagrangian violates the gauge invariance of the theory. To circumvent this we will consider the Stueckelberg extension of the abelian gauge boson in the Weinberg-Salam model stueckelberg ; ruegg . As an explicit example we consider the $Z\rightarrow 2\gamma$ process. $Z$ boson being chargeless is not coupled directly to the photon. We have to look for loop diagrams. Z interacts with the electron and the electron is coupled to the photon. The one loop diagram figure 2 is an example of two photon emission from $Z$ (there will be another diagram with $(p,\nu)\leftrightarrow(k,\lambda)$ which we do not show explicitly). In principle we will have contributions from the muon ($\mu$) and tau ($\tau$) loops as well as quark loops. Their contributions are expected to be small.
@@
$A_{4}~{}=~{}(\vec{s}_{1}\cdot\vec{S})(\vec{s}_{2}\cdot\vec{k})-(\vec{s}_{1}\cdot\vec{k})(\vec{s}_{2}\cdot\vec{S})$ (13)
@@
$T\approx(-1)\int\frac{d^{4}\ell}{(2\pi)^{4}}Tr\left[\gamma^{\mu}\frac{(1-\gamma^{5})}{2}\frac{i}{C_{1}}\gamma^{\lambda}\frac{i}{C_{2}}\gamma^{\nu}\frac{i}{C_{3}}\right]+(p,\nu)\leftrightarrow(k,\lambda)$ (14)
@@
This has to be contracted with the massive gauge boson polarization vectors. As has been pointed out in ruegg , the Stueckelberg extension of the standard model is anomaly free and therefore we can simply estimate the amplitude for the process without worrying about the details of evaluating the loop integral. Including the coupling constants and the masses of the particles involved, we can estimate $|T|^{2}$ (upto numerical factors) to go as
@@
In the above expression $v^{Z}_{e}$ and $v^{A}_{e}$ are, respectively, the effective couplings of the electron with the $Z$ boson and the photon (in the notation of ruegg ), $M_{Z}$ is the mass of the $Z$ boson, $m_{e}$ is the mass of the electron and $m_{\gamma}$ is the mass of the photon. We would like to point out that there are differences between the conventional Weinberg Salam model and the Stueckelberg extension [^1] Here we have projected the $Z-e-e$ vertex to the $V-A$ part since we are only interested in the order of magnitude estimate for the process. However, we note that in the Stueckelberg extension of the standard model there is a slight difference in the electromagnetic coupling of the photon to the left-handed and the right-handed electrons and we refer the reader to the review ruegg for further details.. We see that in the limit $m_{\gamma}\rightarrow 0$ the $Z\rightarrow\gamma\gamma$ does not occur as required by the Landau-Yang theorem.